OpenStudy (anonymous):

Find all real solutions. show work: \[{1 \over (x-1)} + {1 \over (x+2)} = 5/4 \] I get to: \[{2x+1 \over (x-1)(x+2)} = 5/4\] and don't know where to go...

OpenStudy (radar):

I would at that point "cross multiply getting: 5(x-1)(x+2)= 8x + 4 \[5x ^{2}+5x-10=8x+4\]\[5x ^{2}-3x-14=0\] Did you get the same? Now you have a quadratic that can be factored. Can you take it from there?

OpenStudy (radar):

I see you have left. factoring you should get: (x-2)(5x+7)=0

OpenStudy (anonymous):

haven't left. still cramming. lol

OpenStudy (radar):

x=2 x=-7/5

OpenStudy (anonymous):

i actually also got 5x^2 - 3x - 14 = 0 and then stopped. i should have just used the quadratic equation from there! sheesh! thanks!

OpenStudy (radar):

In this case, factoring was simpler than using the quadratic formula, although the formula works just as well.

OpenStudy (anonymous):

i struggle with factoring in this case. lol

OpenStudy (radar):

Yes, it does take some perspective to know if factoring is the way to go. good luck with your studies.